MRM Inc. has additional cash available for investment. One of the production machines needs to be replaced, and management is considering two options. Both options require a similar initial outlay and have a useful life of 10 years. Option #1 will generate $20,000 annually in positive after-tax cash flows and would have an after-tax residual value of $19,000. Option #2 will generate $19,000 annually in positive after-tax cash flows and would have an after-tax residual value of $20,000.

Using a discount rate of 9%, which option is the most attractive?

Select one or more:

a. Option #1

b. Option #2
Answer #1

Present value = Present value of annual cash inflows + present value of residual value

Option 1 = 20000*6.4177 + 19000*0.4224 = $136380

Option 2 = 19000*6.4177 + 20000*0.4224 = $130384

Option #1 is more attractive.